# AR-TTA: A Simple Method for Real-World Continual Test-Time Adaptation

### v1

@@ Abstract
Test-time adaptation is a promising research direction that allows the source model to adapt itself to changes in data distribution without any supervision. Yet, current methods are usually evaluated on benchmarks that are only a simplification of real-world scenarios. Hence, we propose to validate test-time adaptation methods using the recently introduced datasets for autonomous driving, namely CLAD-C and SHIFT. We observe that current test-time adaptation methods struggle to effectively handle varying degrees of domain shift, often resulting in degraded performance that falls below that of the source model. We noticed that the root of the problem lies in the inability to preserve the knowledge of the source model and adapt to dynamically changing, temporally correlated data streams. Therefore, we enhance well-established self-training framework by incorporating a small memory buffer to increase model stability and at the same time perform dynamic adaptation based on the intensity of domain shift. The proposed method, named AR-TTA, outperforms existing approaches on both synthetic and more real-world benchmarks and shows robustness across a variety of TTA scenarios.
@@ 1 Introduction
Deep neural networks have been shown to achieve remarkable performance in various tasks. However, current machine learning models perform very well only when the test-time distribution is close to the training-time distribution. This poses a significant challenge since in a real-world application a domain shift can occur in many circumstances, e.g., weather change, time of day shift, or sensor degradation. For this reason, Test-Time Adaptation (TTA) methods have been widely developed in recent years [37, 35]. Their aim is to adapt the source data pre-trained model to the current data distribution on-the-fly during test-time, using an unlabeled stream of test data. Initial test-time adaptation methods considered a single domain shift at a time. To better simulate real-world challenges, a continual test-time adaptation [39] was recently proposed which involves constantly adapting the model to new domains, which is even more demanding.
Figure 1: Continual test-time adaptation methods evaluated on synthetic (CIFAR-10C) and realistic (CLAD-C) domain shifts. Our method is the only one that consistently allows to improve over the naive strategy of using the (frozen) source model.
@@
An efficient TTA method needs to work well in a wide range of settings. While some domain shifts occur abruptly, there are also several shifts that evolve gradually over time [20]. Additionally, the temporal correlation between consecutive frames violates the i.i.d. assumption. The model should be able to achieve stable performance when handling lengthy sequences, potentially extending indefinitely. Existing approaches are based on updating model parameters using calculated pseudo-labels or entropy regularization [39, 37]. Further, filtering of less reliable samples is often employed to reduce error accumulation and improve the computational efficiency [29, 30, 1]. However, all of the methods can become unstable due to the aforementioned factors, and as a result, the pseudo-labels become noisier and result in performance degradation [5]. Without using any source data the model is prone to catastrophic forgetting [28] of previously acquired knowledge.
@@
Existing methods are mostly evaluated on synthetic datasets or on relatively short-length sequences [37, 39, 6] and as such it is not known how those methods will work in real-life scenarios. Therefore, we adapt the autonomous driving benchmark CLAD [36], to the continual adaptation setting. Moreover, we use a SHIFT dataset [34], which is synthetically generated but is very realistic, provides very long sequences, and allows us to specifically control for different factors (time of day, weather conditions).
@@
In the proposed evaluation setup, we find out that current approaches lack the required stability, as their performance significantly deteriorates compared to the source model, see Figure 1. Additionally, we notice that they struggle to correctly estimate batch norm statistics with temporally correlated data streams and low batch sizes. In our method, we extend popular self-training framework [39, 37] with a small memory buffer, which is used during adaptation to prevent knowledge forgetting, without relying on heuristic-based strategies or resetting model weights that are often used [39, 30]. Thanks to using mixup data augmentation [40] relatively small number of samples are required. Furthermore, we develop a module for dynamic batch norm statistics adaptation, which interpolates the calculated statistics between those of pretrained model and those obtained during deployment, based on the intensity of domain shift. We call our method AR-TTA, as we improve Adaptation by using dynamic batch norm statistics and maintain knowledge by Repeating samples from the memory buffer combined with mixup data augmentation.
As a result, our proposed method AR-TTA is simple, stable, and works well across a range of datasets with different shift intensities, when using small batches of data and over very long sequences. Our main contributions can be summarized as follows:
  • We evaluate and analyze current test-time adaptation methods on realistic, continual domain shift image classification data.
@@
  • We propose a simple continual TTA method based on dynamic batch normalization statistics update and a small memory buffer combined with mixup data augmentation.
@@
  • Extensive evaluation shows that the proposed method obtains state-of-the-art performance on multiple benchmarks with both artificial distortions and real-life ones from autonomous driving.
@@ 2 Related Work
Test-time adaptation (TTA). Domain adaptation methods can be split into different categories based on what information is assumed to be available during adaptation [38]. While in some scenarios, access to some labels in target distribution is available, the most common is unsupervised domain adaptation, which assumes that the model has access to labeled source data and unlabeled target data at adaptation time. Popular approaches are based on either minimizing the discrepancy between the source and target feature distributions [12, 17]. Alternative approaches are based on self-training, which uses the model’s predictions on the target domain as pseudo-labels to guide the model adaptation [42, 22].
Additionally, in test-time adaptation the model needs to adapt to the test-time distribution on the fly, in an online fashion. In the test-time training (TTT) method [35] the model solves self-supervised tasks on the incoming batches of data to update its parameters. TENT [37] updates only batch-norm statistics to minimize predictions entropy. This assumes that updating only batch norm statistics is sufficient to solve the problem, which might not be the case for real-world scenarios. EATA [29] further improves the efficiency of test-time adaptation methods, by using only diverse and reliable samples (with low prediction entropy). Additionally, it uses EWC [18] regularization to prevent drastic changes in parameters important for the source domain.
Contrary to the TENT and EATA approaches, CoTTA [39] updates the whole model. To prevent performance degradation it uses exponential weight averaging as well as stochastic model restoration, where randomly selected weights are reset to the source model. SAR [30] further improves by removing noisy test samples with large gradients and adding loss components that encourage model weights to go to a flat minimum. Nevertheless, they also use model reset, to prevent forgetting.
TTA benchmarks. The most popular setting for test-time adaptation includes using different classes of synthetic corruptions proposed in [14], which are then utilized for test-time adaptation, one at a time (so allowing model reset between different domains). However, in practical applications, the target distribution can easily change perpetually over time, e.g., due to changing weather and lightness conditions, or due to sensor corruptions, hence the setting of continual test-time adaptation was recently introduced [39]. The authors proposed to use a continual version of the corrupted benchmark (that is without model reset at domain boundaries).
@@
Another popular dataset for continual test-time adaptation is DomainNet [31] which consists of images in different domains (e.g., sketches, infographics). Yet, the distribution shifts arising in the real world may be very different from the synthetic ones. Hence, recently a CLAD, autonomous driving benchmark [36] was introduced. It consists of naturally occurring distribution shifts like changes in weather and lighting conditions, traffic intensity, etc. It was developed for the supervised Continual Learning scenario. In this work, we use it for test-time adaptation, that is without using any label information. In our work, we also use SHIFT benchmark [34], a synthetically generated dataset for autonomous driving with realistic discrete and continuous shifts. Similarly to use, CoTTA [39] includes realistic domain shifts but their test is very small (1600 images). To sum up, we extend over previous TTA work by focusing on realistic continual domain shifts over very long sequences.
Continual Learning. Our work is also inspired by continual learning where the learner is presented with data from different tasks in a sequential fashion. Without access to the previous data from previous tasks, the model is prone to catastrophic forgetting [28]. Popular approaches use knowledge distillation to regularize changes in outputs of the model (compared to the model train on source task) [21] or regularize changes in model parameters [18]. On the other hand, exemplar-based approaches assume some limited access to the source data, which greatly helps to reduce the forgetting [27]. Further, some continual learning approaches focus on online setting, where each data point can be visited only once and small batch sizes are assumed [2].
@@ 3 Method
Figure 2: Our method AR-TTA adapts to the domain of an unlabeled continual stream of data utilizing exemplars of source pre-training data saved in the memory. We keep two identical neural network models with different sets of weights: the teacher model and the student model. We sample one exemplar from memory for each chosen image and mixup image pairs. Similarly, pseudo-labels from the teacher model are mixupped with the labels sampled from memory exemplars. The student model is updated based on cross-entropy loss between its predictions on augmented samples and augmented pseudo-labels. The teacher model is adapted based on an exponential moving average of student’s weights. Predictions for each image are taken from the teacher model.
The aim of TTA is to adapt the pre-trained model $f_{\theta_{0}}$ trained on the labeled source data $(\mathcal{X}^{S},\mathcal{Y}^{S})$ to the ever-changing stream of unlabeled test data batches $\textbf{x}^{T}$ on the fly during the evaluation.
Our proposed approach to TTA (AR-TTA) can be divided into three parts. We start the description by introducing the model update procedure in Subsection 3.1. Then, in Subsection 3.2 we explain the usage of experience replay with mixup augmentation. The process of adapting batch normalization statistics is presented in Subsection 3.3 The overview of our method is presented in Figure 2.
@@ 3.1 Weight-averaged Consistency
Updating the model’s weights during test-time is not a trivial task, considering the lack of data labels and the possibility of error accumulation due to noisy training feedback. Moreover, with the update procedure comes the stability-plasticity dilemma. The model should be stable enough to minimize the risk of deteriorating performance, model collapse, and catastrophic forgetting [11]. On the other hand, we want the model to be flexible to keep up with domain changes and adapt on time. Following previous works [9, 39], which show effective methods for alleviating mentioned challenges, we propose to employ self-training on pseudo-labels and keep two models, where one is updated by the exponential moving average of another’s weights.
We initialize two identical artificial neural network models, student model $f_{\theta}$ and teacher model $f_{\theta^{\prime}}$, with the identical weights obtained by training on source data. For each batch of test data $\textbf{x}_{t}^{T}$ at time step $t$ we generate predictions from both models. Teacher model predictions $\hat{y^{{}^{\prime}}}_{t}^{T}$ are used as soft pseudo-labels. The student model is updated by the cross-entropy loss between its predictions and the pseudo-labels:
$\mathcal{L}_{\theta_{t}}(\textbf{x}_{t}^{T})=-\sum_{c}\hat{y^{{}^{\prime}}}_{t,c}^{T}\log\hat{y}_{t,c}^{T}$ (1)
where $\hat{y}_{t,c}^{T}$ is the probability of class $c$ predicted by the student model.
Next, teacher’s weights $\theta^{{}^{\prime}}$ are updated by exponential moving average of student’s weights $\theta$:
$\theta_{t+1}^{{}^{\prime}}=\alpha\theta_{t}^{{}^{\prime}}+(1-\alpha)\theta_{t+1}$ (2)
where $\alpha$ is a smoothing factor.
As mentioned in  [39] using a weight-averaged teacher model ensures less noisy pseudo-labels since weight-averaged models over training steps often yield more accurate predictions than the final model and the added inertia prevents hasty, rapid weights update based on noisy self-training feedback. Moreover, susceptibility to catastrophic forgetting is decreased considering the fact that the weights are the combination of past iterations.
We do not limit the weights update only to affine parameters of batch normalization layers, as in many other TTA methods [29, 37, 30], and we update the whole model. We argue that adopting only batch normalization layers does not give the model enough flexibility to perform successfully on varying domains. We confirm this claim experimentally in the appendix and show that updating the whole models provides the best results, compared to fine-tuning only batch-norm layers or different blocks of the model.
The final predictions for the current test batch $\textbf{x}_{t}^{T}$ are the classes with the highest probabilities in pseudo-labels generated by the teacher model before the update.
@@ 3.2 Experience Replay with Adaptation
During continual test-time adaptation to unlabeled data, the model is exposed to training feedback that most likely differs from what it has learned during source pre-training. Pseudo-labels or entropy minimization feedback are not guaranteed to be accurate, and frequent model updates inevitably strive for significant error accumulation. These factors can be the cause of the model forgetting the initial knowledge. Noisy self-training and significant forgetting can even cause the model to collapse, as shown in [30]. For practical applications, a utilized method has to be reliable and the risk of collapsing has to be reduced to a minimum.
To alleviate this issue, we propose to use the class-balanced replay buffer of exemplars during adaptation to remind the model what it has learned and strengthen its initial knowledge. We take inspiration from continual learning approaches, which show that exemplars are one of the most effective approaches for this task [32, 4, 3, 36]. To fully take advantage of exemplars and make the latent representations of a model more robust for a given task, we follow a few of the continual learning works [24, 41] and propose to use Mixup data augmentation [40].
After completing the pre-training of the source model, we store a specific number of random exemplars from the labeled source data in the memory. The number of exemplars is the same for all classes. This works better than a random selection of exemplars, which we show in the appendix. In each test-time adaptation iteration, we randomly sample exemplars $\textbf{x}_{t}^{S}$, along with their labels $\textbf{y}_{t}^{S}$, from memory. The number of sampled exemplars is equal to the batch size. Mixupped batch of samples $\tilde{\textbf{x}}_{t}$ is generated by linearly interpolating samples from test data with samples from memory:
@@
where $\lambda\sim$ Beta$(\psi,\rho)$, for $\psi,\rho\in(0,\infty)$. Similarly, labels for cross-entropy loss are the result of interpolation between pseudo-labels produced by the teacher model based on the current unmodified test batch $\hat{y}_{t}^{\prime T}$ and labels $\textbf{y}_{t}^{S}$ from the memory, with the same $\lambda$ parameter value:
$\tilde{\textbf{y}}_{t}=\lambda\hat{\textbf{y}^{{}^{\prime}}}_{t}^{T}+(1-\lambda)\textbf{y}_{t}^{S}$ (4)
Student model takes augmented batch $\tilde{\textbf{x}}_{t}$ as input. Its predictions are compared with interpolated labels $\tilde{\textbf{y}}_{t}$ to calculate the loss as described in the previous Subsection 3.1.
A similar approach to mixing exemplars from replay memory with the ones to train on was successfully used in LUMP [24], however, they used this method for the continual learning tasks.
Using experience replay along with the Mixup augmentation, helps the model preserve already obtained knowledge. Furthermore, having pseudo-labels mixed with certainly accurate labels for exemplars, makes the noisiness of pseudo-labels less impactful for the adaptation process.

### v2

@@ Abstract
Test-time adaptation is a promising research direction that allows the source model to adapt itself to changes in data distribution without any supervision. Yet, current methods are usually evaluated on benchmarks that are only a simplification of real-world scenarios. Hence, we propose to validate test-time adaptation methods using the recently introduced datasets for autonomous driving, namely CLAD-C and SHIFT. We observe that current test-time adaptation methods struggle to effectively handle varying degrees of domain shift, often resulting in degraded performance that falls below that of the source model. We noticed that the root of the problem lies in the inability to preserve the knowledge of the source model and adapt to dynamically changing, temporally correlated data streams. Therefore, we enhance the well-established self-training framework by incorporating a small memory buffer to increase model stability and at the same time perform dynamic adaptation based on the intensity of domain shift. The proposed method, named AR-TTA, outperforms existing approaches on both synthetic and more real-world benchmarks and shows robustness across a variety of TTA scenarios. The code is available at https://github.com/dmn-sjk/AR-TTA.
@@ 1 Introduction
[…]
@@
An effective TTA method should work well regardless of a wide range of challenges, encompassing both abrupt and gradual domain shifts. Moreover, the model should be able to maintain stable performance even when handling lengthy sequences, potentially extending indefinitely. Existing approaches are based on updating model parameters using pseudo-labels or entropy regularization [Wang et al.(2022)Wang, Fink, Gool, and Dai, Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell]. Further, filtering the less reliable samples is often employed to reduce error accumulation and improve the computational efficiency [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan, Alfarra et al.(2023)Alfarra, Itani, Pardo, Alhuwaider, Ramazanova, Pérez, Cai, Müller, and Ghanem]. However, all of the methods can become unstable due to the aforementioned difficulties, accumulate errors, make the pseudo-labels noisier, and cause performance degradation [Chen et al.(2019)Chen, Xie, Huang, Rong, Ding, Huang, Xu, and Huang]. Without using any source data the model is prone to catastrophic forgetting [McCloskey and Cohen(1989)] of previously acquired knowledge.
Figure 1: Continual test-time adaptation methods evaluated on artificial (CIFAR10C, ImageNet-C [Hendrycks and Dietterich(2019)]) and natural (CIFAR10.1 [Recht et al.(2018b)Recht, Roelofs, Schmidt, and Shankar], SHIFT [Sun et al.(2022)Sun, Segu, Postels, Wang, Van Gool, Schiele, Tombari, and Yu], CLAD-C [Verwimp et al.(2023)Verwimp, Yang, Parisot, Hong, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars]) domain shifts. Our method is the only one that consistently allows to improve over the naive strategy of using the (frozen) source model.
@@
Currently, methods are mostly evaluated on datasets with artificial, unrealistic domain shifts or relatively short-length sequences [Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell, Wang et al.(2022)Wang, Fink, Gool, and Dai, Chen et al.(2022)Chen, Wang, Darrell, and Ebrahimi], hence it is unknown how those methods will work in real-life scenarios with natural domain shifts. Therefore, we adapt the autonomous driving benchmark CLAD [Verwimp et al.(2023)Verwimp, Yang, Parisot, Hong, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars], to the continual adaptation setting. Moreover, we use a SHIFT dataset [Sun et al.(2022)Sun, Segu, Postels, Wang, Van Gool, Schiele, Tombari, and Yu], which is synthetically generated by utilizing the realistic autonomous driving simulator CARLA [Dosovitskiy et al.(2017)Dosovitskiy, Ros, Codevilla, Lopez, and Koltun], provides very long sequences, and allows us to specifically control for different domain factors (time of day, weather conditions).
@@
In the proposed evaluation setup, we find out that current approaches lack the required stability, as their performance significantly deteriorates compared to the source model (Fig. 1). Additionally, we notice that they struggle to correctly estimate batch norm statistics with temporally correlated data streams and low batch sizes. In our TTA method for an image classification task, we extend popular self-training framework [Wang et al.(2022)Wang, Fink, Gool, and Dai] with a small memory buffer, which is used during adaptation to prevent knowledge forgetting, without relying on heuristic-based strategies or resetting model weights that are often used [Wang et al.(2022)Wang, Fink, Gool, and Dai, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. Thanks to using mixup data augmentation [Zhang et al.(2018)Zhang, Cisse, Dauphin, and Lopez-Paz], a relatively small number of samples is required. Furthermore, we develop a module for dynamic batch norm statistics adaptation, which interpolates the calculated statistics of the pre-trained model and those obtained during testing, based on the intensity of domain shift. We call our method AR-TTA, as we improve Adaptation by using dynamic batch norm statistics and maintain knowledge by Repeating samples from the memory buffer combined with mixup data augmentation. As a result, our proposed method AR-TTA is simple, stable, and works well across a range of datasets with different shift intensities, when using small batches of data and over very long sequences. Our main contributions can be summarized as follows:
  • We propose novel evaluation benchmarks for TTA based on autonomous-driving scenarios, which show significant limitations of existing TTA methods.
@@
  • Based on discovered vulnerabilities, we propose a TTA method that dynamically updates the batch norm statistics based on the intensity of domain shift and use small memory buffer combined with mixup data augmentation.
@@
  • Extensive evaluation shows that the proposed method obtains state-of-the-art performance on benchmarks with both artificial distortions and natural domain shifts.
@@ 2 Related Work
Test-time adaptation (TTA). Domain adaptation methods can be split into different categories based on what information is assumed to be available during adaptation [Wang and Deng(2018)]. While in some scenarios, access to some labels in target distribution is available, the most common scenario is unsupervised domain adaptation, in which the model has access to labeled source data and unlabeled target data at adaptation time. Additionally, in test-time adaptation the model needs to adapt to the test-time distribution on-the-fly, in an online fashion. In the test-time training (TTT) method [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt] the model solves self-supervised tasks on the incoming batches of data to update its parameters. TENT [Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell] updates only batch-norm weights to minimize predictions entropy. EATA [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan] further improves the efficiency of test-time adaptation methods, by using only diverse and reliable samples (with low prediction entropy). Additionally, it uses EWC [Kirkpatrick et al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness, Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, et al.] regularization to prevent drastic changes in parameters important for the source domain. Contrary to the TENT and EATA approaches, CoTTA [Wang et al.(2022)Wang, Fink, Gool, and Dai] updates the whole model. To prevent performance degradation it uses exponential weight averaging as well as stochastic model restoration, where randomly selected weights are reset to the source model. SAR [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan] further improves by removing noisy test samples with large gradients and using sharpness-aware minimization. Nevertheless, they also use model reset, to prevent forgetting. RMT [Döbler et al.(2022)Döbler, Marsden, and Yang] propose a mean teach framework and adapt it based on contrastive loss, self-training loss, and a simple replay strategy utilizing all of the source training data during test-time. Moreover, it assumes an additional warm-up step before the adaptation. Our method uses a small number of replay exemplars and does not require any extra steps before TTA.
@@
TTA benchmarks. The most popular setting for test-time adaptation includes applying different classes of synthetic corruptions proposed in [Hendrycks and Dietterich(2019)] to commonly used datasets, like CIFAR10 [Krizhevsky(2009)] or ImageNet [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei] to artificially obtain domain shifts. Another popular dataset for continual test-time adaptation is DomainNet [Peng et al.(2019)Peng, Bai, Xia, Huang, Saenko, and Wang] which consists of images in different domains (e.g., sketches, infographics). Yet, the distribution shifts arising in real-world deployment, e.g., in autonomous driving, may be very dissimilar from different renditions of the same objects. Hence, recently a CLAD, autonomous driving benchmark [Verwimp et al.(2023)Verwimp, Yang, Parisot, Hong, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars] was introduced. It consists of naturally occurring distribution shifts like changes in weather and lighting conditions, traffic intensity, etc. It was developed for the supervised continual learning scenario. In this work, we use it for test-time adaptation, that is without using any label information. In our work, we also use SHIFT benchmark [Sun et al.(2022)Sun, Segu, Postels, Wang, Van Gool, Schiele, Tombari, and Yu], a synthetically generated dataset for autonomous driving that captures the continuously evolving nature of the real world. Similarly to us, CoTTA [Wang et al.(2022)Wang, Fink, Gool, and Dai] includes realistic domain shifts but their test is very small (1600 images). To sum up, we extend over previous TTA work by focusing on realistic continual domain shifts over very long sequences.
@@ 3 Method
Figure 2: Our method, AR-TTA, utilizes a replay strategy and the mean teacher framework. Each image is paired with an exemplar sampled from memory and image pairs are mixed up. Similarly, pseudo-labels from the teacher model are mixed up with the labels of sampled exemplars. The student model is updated based on cross-entropy loss between its predictions on augmented samples and augmented pseudo-labels. The teacher model is adapted based on an exponential moving average of student’s weights. Predictions for each image are taken from the teacher model.
Continual TTA aims to adapt the pre-trained model $f_{\theta_{0}}$ trained on the labeled source data $(\mathcal{X}^{S},\mathcal{Y}^{S})$ to the ever-changing stream of unlabeled test data batches $\textbf{x}^{T}$ on-the-fly during the evaluation. Our proposed approach (Fig. 2.) can be divided into three parts. We start the description by introducing the model update procedure in Section 3.1. Then, in Section 3.2 we explain the usage of experience replay with mixup augmentation. The process of adapting batch normalization statistics is presented in Section 3.3
@@ 3.1 Weight-averaged Consistency
Following previous works [Döbler et al.(2022)Döbler, Marsden, and Yang, Wang et al.(2022)Wang, Fink, Gool, and Dai], we use the self-training mean teacher framework to update the model. It improves the reliability of pseudo-labels and minimizes the risk of error accumulation. We initialize two identical artificial neural network models, student model $f_{\theta}$ and teacher model $f_{\theta^{\prime}}$, with the equal weights obtained by training on source data. For each batch of test data $\textbf{x}_{t}^{T}$ at time step $t$, predictions are derived from both models. Teacher model predictions $\hat{y}_{t}^{\prime T}$ are used as soft pseudo-labels. The student model is updated by the cross-entropy loss between its predictions and the pseudo-labels:
$\mathcal{L}_{\theta_{t}}(\textbf{x}_{t}^{T})=-\sum_{c}\hat{y}_{t,c}^{\prime T}\log\hat{y}_{t,c}^{T}$ (1)
where $\hat{y}_{t,c}^{T}$ is the probability of class $c$ predicted by the student model. Next, teacher’s weights $\theta^{\prime}$ are updated by exponential moving average of student’s weights $\theta$:
$\theta_{t+1}^{\prime}=\alpha\theta_{t}^{\prime}+(1-\alpha)\theta_{t+1}$ (2)
where $\alpha$ is a smoothing factor. Using a weight-averaged teacher model ensures less noisy pseudo-labels [Wang et al.(2022)Wang, Fink, Gool, and Dai], improves generatlization [Morales-Brotons et al.(2024)Morales-Brotons, Vogels, and Hendrikx], and added inertia prevents rapid weights update based on noisy self-training feedback.
We do not limit the weights update only to affine parameters of batch normalization layers, as in many other TTA methods [Niu et al.(2022)Niu, Wu, Zhang, Chen, Zheng, Zhao, and Tan, Wang et al.(2021)Wang, Shelhamer, Liu, Olshausen, and Darrell, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan], and we adapt the whole model. We argue that adapting only batch normalization layer weights does not give the model enough flexibility to perform successfully on varying domains. We confirm this claim experimentally in the appendix.
@@ 3.2 Experience Replay with Adaptation
During continual test-time adaptation, self-training feedback is not guaranteed to be accurate, and frequent model updates inevitably strive for significant error accumulation, catastrophic forgetting, and even model collapse [Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. Continual learning works show that experience replay is one of the most effective strategies for mitigating catastrophic forgetting [Rebuffi et al.(2017)Rebuffi, Kolesnikov, Sperl, and Lampert, Chaudhry et al.(2019)Chaudhry, Gordo, Dokania, Torr, and Lopez-Paz, Boschini et al.(2023)Boschini, Bonicelli, Buzzega, Porrello, and Calderara, Verwimp et al.(2023)Verwimp, Yang, Parisot, Hong, McDonagh, Pérez-Pellitero, De Lange, and Tuytelaars]. Following this insight we propose to use the replay strategy during adaptation to remind the model what it has learned and strengthen its initial knowledge. Building on this, we integrate Mixup data augmentation [Zhang et al.(2018)Zhang, Cisse, Dauphin, and Lopez-Paz] to enhance model robustness, inspired by works such as [Madaan et al.(2022)Madaan, Yoon, Li, Liu, and Hwang, Zhu et al.(2021)Zhu, Cheng, Zhang, and Liu].
After completing the pre-training of the source model, we store a low, equal for each class, number of random exemplars from the labeled source data in the memory. In each test-time adaptation iteration, we randomly sample exemplars $\textbf{x}_{t}^{S}$, along with their labels $\textbf{y}_{t}^{S}$, from memory. The number of sampled exemplars is equal to the batch size. Mixed up batch of samples $\tilde{\textbf{x}}_{t}$ is generated by linearly interpolating samples from test data with samples from memory:
@@
where $\lambda\sim$ Beta$(\psi,\rho)$, for $\psi,\rho\in(0,\infty)$. Similarly, labels for cross-entropy loss are the result of interpolation between pseudo-labels produced by the teacher model based on the current unmodified test batch $\hat{y}_{t}^{\prime T}$ and labels $\textbf{y}_{t}^{S}$ from the memory, with the same $\lambda$:
$\tilde{\textbf{y}}_{t}=\lambda\hat{\textbf{y}}_{t}^{\prime T}+(1-\lambda)\textbf{y}_{t}^{S}$ (4)
Student model takes augmented batch $\tilde{\textbf{x}}_{t}$ as input. Its predictions are compared with interpolated labels $\tilde{\textbf{y}}_{t}$ to calculate the loss as described in the previous Section 3.1.
A similar approach to mixing exemplars from replay memory with the ones to train on was successfully used in LUMP [Madaan et al.(2022)Madaan, Yoon, Li, Liu, and Hwang], however, they used this method for the continual representation learning task. Using experience replay along with the Mixup augmentation helps the model preserve already obtained knowledge and makes the noisiness of pseudo-labels less impactful for the adaptation process.